how much pure acid should be added to 18 L of 30% acid to increase the concentration of 50% acid?
Answer by stanbon(75887)

how much pure acid should be added to 18 L of 30% acid to increase the concentration of 50% acid?
-----
Equation:
acid + acid = acid
0.30*18 + 1*x = 0.50(18+x)
----
Multiply thru by 100 to get:
30*18 + 100x = 50*18 + 50x
50x = 20*18
x = (2/5)(18)
x = 36/5 = 7 1/5 L (amt of pure acid needed)
==============
Cheers,
Stan H.